# secReferences

### v1

@@ Appendix A Proofs
We encode all the exchange plans in $\mathcal{F}$ based on the vertex inclusion. Let =$\mathcal{F}=\{F_{1},...,F_{S}\}$, where $S=|\mathcal{F}|$, and $F_{i}\in\mathbb{R}^{N}$ where $F_{ij}$ is the binary indicator that the $j$th vertex is included in the exchange plan $F_{i}$. Let $\{w_{1},...w_{S}\}$ be the total utilities of $\{F_{1},...,F_{S}\}$, and let $\{p_{1},...,p_{S}\}$ be the assigned probabilities of $\{F_{1},...,F_{S}\}$. We write $F=(F_{ij})_{S\times N}$, $p=(p_{1},...,p_{S})^{\mathrm{\scriptscriptstyle T}}$, and $w=(w_{1},...,w_{S})^{\mathrm{\scriptscriptstyle T}}$. Then, $q=F^{\mathrm{\scriptscriptstyle T}}p$ is the vector of probability of being matched for each vertex, and $w^{\mathrm{\scriptscriptstyle T}}p$ is the expected utility. Let $\overline{q}_{ij}=|V_{ij}|^{-1}\sum_{v\in V_{ij}}q_{v}$ denote the average selection rate in $V_{ij}$. Then, we can write $(\overline{q}_{11}-\overline{q}_{01},...,\overline{q}_{1M}-\overline{q}_{0M})^{\mathrm{\scriptscriptstyle T}}=Z^{\mathrm{\scriptscriptstyle T}}p$ for some known $S\times M$ matrix $Z$. Let $z_{i}\in\mathbb{R}^{M}$ denote the $i$th row of $Z$, and let $\tilde{z}_{i}=(w_{i}\ z_{i}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{M+1}$. Notice that $Z^{\mathrm{\scriptscriptstyle T}}p$ is a convex combination of $z_{1},...,z_{S}$. We consider $\Omega$ to be the convex hull of $\{\tilde{z}_{1},...,\tilde{z}_{S}\}$. Let $\Delta=[-l_{1},l_{1}]\times...\times[-l_{M},l_{M}]$.
Then, the optimization problem (4) searches over $\Omega\cap(\mathbb{R}\times\Delta)$ and returns an optimum point with the maximum first coordinate. Since $S$ is finite, the convex hull $\Omega$ is compact in $\mathbb{R}^{M+1}$. Thus, $\Omega\cap(\mathbb{R}\times\Delta)$ is compact in $\mathbb{R}^{M+1}$. Furthermore, the optimum point is on the boundary of $\Omega\cap(\mathbb{R}\times\Delta)$, which is also on the boundary of $\Omega$. By Lemma S1, the optimum point can be represented as a convex combination of $\tilde{z}_{1},...,\tilde{z}_{S}$ with at most $M+1$ nonzero coefficients. ∎
The proof of Proposition 2 depends on the following lemma.
Lemma S2.
Consider a random $k$-partite graph $(A_{1},A_{2},...,A_{k})$, where $A_{1},...,A_{k}$ contain $n$ vertices. Each possible edges appears independently with probability no less than $p>0$. Then, the $k$-partite graph has a perfect matching almost surely.
Proof of Lemma S2.
When $k=2$, Erdős and Rényi (1968) gives the probability of perfect matching, which immediately implies Lemma S2. When $k>2$, it can be shown by mathematical induction. We omit the details. ∎
@@
We use $(b_{1},b_{2},r,a)$ to denote the vertex type of an incompatible donor-patient pair with donor blood type $b_{1}$, patient blood type $b_{2}$, patient sensitization level $r$, and patient sensitive group level $a$. Let $\mathbb{T}=\{O,A,B,AB\}\times\{O,A,B,AB\}\times\{r_{1},...,r_{M}\}\times\{0,1\}$. Then, we partition the cycle or relevant subset set $\mathcal{C}$ based on the graph isomorphism with respect to the vertex types in $\mathbb{T}$. Specifically, for $c_{1},c_{2}\in\mathcal{C}$, we view cycle or relevant subset $c_{1}$ and $c_{2}$ of the same type, if there exists a one-to-one mapping $\phi$ from the vertex set in $c_{1}$ to the vertex set in $c_{2}$, satisfying that any $v$ and $\phi(v)$ are of the same type and $(v_{1},v_{2})\in E$ if and only if $(\phi(v_{1}),\phi(v_{2}))\in E$. Therefore, if $c_{1},c_{2}$ are of the same type, then $u(c_{1})=u(c_{2})$. Suppose there are $K$ such types of cycles or relevant subsets in $\mathcal{C}$. Let $U\in\mathbb{R}^{K}$, where $U_{i}=u(c)$ for any cycle or relevant subset $c$ of the $i$th type in $\mathcal{C}$. For arbitrary graph $G$, we encode any exchange allocation as a vector $Z\in\mathbb{R}^{K}$, where $Z_{i}$ denotes the number of cycle or relevant subgraph of the $i$th type in the exchange plan. Let $Opt(G)$ be the solution of optimization problem (7). We claim that the following result holds almost surely as $N\to\infty$, which implies the result in Proposition 2.
  1. 1.
    There exists $Y\in\mathbb{R}^{K}$ such that $Z=\lfloor NY\rfloor\in\mathbb{R}^{K}$ is an achievable exchange allocation in $G$, and $Opt(G)-U^{{\mathrm{\scriptscriptstyle T}}}Z$ is $o(N)$.
  2. 2.
    We can choose $Y\in\mathbb{R}^{K}$ such that $Z=\lfloor NY\rfloor\in\mathbb{R}^{K}$ is an achievable exchange allocation in $G$, where the subgroup $V_{b_{1},b_{2},i,j}$ is prioritized over $V_{b_{1},b_{2},1-i,j}$ for all $(b_{1},b_{2},i,j)\in\mathcal{P}$.
[…]
We can prove the above claims below. Let $(\Omega,\mathcal{F},P)$ denote the probability space. We use $G(N)$ to denote the random graph of size $N$, and we use $G(N;\omega)$ to denote the realized graph of size $N$ for some $\omega\in\Omega$. For fixed $N$, the set $\{Opt(G(N;\omega)):\omega\in\Omega\}$ is bounded, because the number of the edges in $G(N)$ and the edge utilities are finite. Let $S_{N}$ denote the supremum of the set $\{Opt(G(N;\omega)):\omega\in\Omega\}$. Let $Z(N;\omega)\in\mathbb{R}^{K}$ denote the optimal exchange allocation in $G(N;\omega)$. Then, Bolzano–Weierstrass Theorem implies that there exists a subsequence in $\{z(N;\omega):\omega\in\Omega\}$ that converges to $Z_{N}\in\mathbb{R}^{K}$ elementwise, satisfying $U^{{\mathrm{\scriptscriptstyle T}}}Z_{N}=S_{N}$.
For diverging $N$, the sequence $\{S_{N}/N:N\in\mathbb{N}\}$ is bounded, so there exists a subsequence in $\{S_{N}/N:N\in\mathbb{N}\}$ that converges to the supremum of $\{S_{N}/N:N\in\mathbb{N}\}$. Thus, there exists a sequence of increasing numbers $i_{1}<i_{2}<i_{3}<...$ in $\mathbb{N}$, such that $\{S_{i_{k}}/{i_{k}}:k\in\mathbb{N}\}$ converges to the supremum of $\{S_{N}/N:N\in\mathbb{N}\}$. Since the set $\{Z_{i_{k}}/i_{k}:k\in\mathbb{N}\}$ is bounded elementwise, Bolzano–Weierstrass Theorem implies that there exists a subsequence in $\{Z_{i_{k}}/i_{k}:k\in\mathbb{N}\}$ that converges to $Y\in\mathbb{R}^{K}$ elementwise. Based on the construction, we have that $S_{N}/(NU^{{\mathrm{\scriptscriptstyle T}}}Y)$ converges to 1, and thus the supremum of $\{Opt(G(N;\omega))/(NU^{{\mathrm{\scriptscriptstyle T}}}Y):\omega\in\Omega\}$ is no greater than 1.
Next, we show that $Z=\lfloor NY\rfloor$ is an achievable exchange allocation in $G$ almost surely as $N\to\infty$, which will further implies that $Opt(G)-U^{{\mathrm{\scriptscriptstyle T}}}Z$ is $o(N)$. We randomly divide the vertices of $G$ into $K$ disjoint subgraphs based on allocation $Z$, satisfying that the $i$th subgraph contains sufficient vertices to match $Z_{i}$ cycles or relevant subgraphs of the $i$th type. Since some elements in $Z$ can be zero, these corresponding subgraphs can be empty. Lemma S2 guarantees that there is a perfect matching in every subgraph almost surely. Therefore, $Z$ is achievable in random graph $G$ almost surely as $N\to\infty$.
We can manually adjust the elements in $Y$ such that the subgroup $V_{b_{1},b_{2},i,j}$ is prioritized over $V_{b_{1},b_{2},1-i,j}$ for all $(b_{1},b_{2},i,j)\in\mathcal{P}$. Similar to the above argument, $Z=\lfloor NY\rfloor$ is an achievable exchange allocation in $G$ almost surely as $N\to\infty$. We have finished the proof of the above two claims. ∎
The following proofs are based on the expected proportions of subgroups. The observed size of subgroups divided by $N$, should converge to the corresponding expected proportions of subgroups, with high probability. These differences are ignorable in the asymptotic upper bounds of the price of fairness.

### v2

@@ Appendix A Proofs
We organize the proof into six steps:
  • Step 1 defines coarse and fine types and their frequencies.
  • Step 2 introduces template types and a coarse-type linear program.
  • Step 3 shows that the unconstrained optimal value is asymptotic to the LP optimum.
  • Step 4 splits coarse usage into fine types in a way that encodes priorities.
  • Step 5 realizes these fine-type counts in $G_{N}$ using Lemma S1.
  • Step 6 combines everything.
Throughout, we work under the assumptions of Section 4.
Step 1: Coarse and fine types and their frequencies. Each vertex $v$ (an incompatible donor–patient pair) has
$b_{1}(v),b_{2}(v)\in\{O,A,B,AB\},\quad r(v)\in\{r_{1},\dots,r_{M}\},\quad A(v)\in\{0,1\}.$
Define the coarse type
$\tilde{\tau}(v):=(b_{1}(v),b_{2}(v),r(v))\in\widetilde{\mathbb{T}}:=\{O,A,B,AB\}^{2}\times\{r_{1},\dots,r_{M}\},$
and the fine type
$\tau(v):=(\tilde{\tau}(v),A(v))\in\mathbb{T}:=\widetilde{\mathbb{T}}\times\{0,1\}.$
For each $N$, let $V_{N}$ denote the vertex set of $G_{N}$, with $|V_{N}|=N$. For $t\in\mathbb{T}$ and $\tilde{t}\in\widetilde{\mathbb{T}}$, define
$n_{t}(N):=\bigl|\{v\in V_{N}:\tau(v)=t\}\bigr|,\qquad n_{\tilde{t}}(N):=\sum_{a=0}^{1}n_{(\tilde{t},a)}(N),$
and let $\pi_{t}>0,\pi_{\tilde{t}}>0$ be the corresponding type probabilities under the i.i.d. type distribution.
By the strong law of large numbers, almost surely
$\frac{n_{t}(N)}{N}\to\pi_{t},\qquad\frac{n_{\tilde{t}}(N)}{N}\to\pi_{\tilde{t}}\quad\text{as }N\to\infty\ \text{for all }t,\tilde{t}$. (S1)
We henceforth work on any realization for which (S1) holds.
Step 2: Template types and a coarse-type linear program. Ignoring the protected feature, we classify cycles (or relevant subsets) into finitely many “templates”. Two sets $c_{1},c_{2}\in\mathcal{C}(G_{N})$ have the same template type if there is a bijection $\phi:V(c_{1})\to V(c_{2})$ such that:
  • $\tilde{\tau}(v)=\tilde{\tau}(\phi(v))$ for all $v\in V(c_{1})$; and
  • $(v_{1},v_{2})\in E_{N}$ iff $(\phi(v_{1}),\phi(v_{2}))\in E_{N}$.
Because the size $|c|$ is uniformly bounded and $\widetilde{\mathbb{T}}$ is finite, there are only finitely many template types. Index them by $i=1,\dots,K$.
For each template $i$, fix a representative directed pattern $H_{i}$ with vertex set $V(H_{i})$ and $|V(H_{i})|=s_{i}$. For each coarse type $\tilde{t}\in\widetilde{\mathbb{T}}$, let
[…]
Since $u(c)$ depends only on the induced subgraph and the coarse types (and is invariant under relabelings preserving this information), $u_{i}:=u(c)$ is well-defined for each template $i$.
We consider the coarse-type linear program
$\max_{z\geq 0}$ $U^{\mathrm{\scriptscriptstyle T}}z=\sum_{i=1}^{K}u_{i}z_{i},$ (S2)
s.t. $\sum_{i=1}^{K}\tilde{a}_{\tilde{t},i}z_{i}\leq\pi_{\tilde{t}},\quad\forall\tilde{t}\in\widetilde{\mathbb{T}}$.
Feasibility is clear and boundedness follows from $u(c)/|c|\leq C$ and bounded template sizes. Let $\rho$ be the optimal value and $z^{\star}$ an optimal solution.
Step 3: Asymptotic behavior of the unconstrained optimum. We show that
$\mathrm{Opt}(G_{N})=\rho N+o(N)\quad\text{almost surely.}$ (S3)
Step 3A: Upper bound. Let $(x_{c})_{c\in\mathcal{C}(G_{N})}$ be an optimal solution of the unconstrained problem (5) on $G_{N}$. For each template type $i$, define
$Z^{(N)}_{i}:=\sum_{c\in\mathcal{C}(G_{N})}x_{c}\,1_{\{c\text{ has template type }i\}},\qquad y^{(N)}:=\frac{1}{N}Z^{(N)}.$
Vertex-disjointness implies that, for each coarse type $\tilde{t}$,
$\sum_{i=1}^{K}\tilde{a}_{\tilde{t},i}Z^{(N)}_{i}\leq n_{\tilde{t}}(N).$
Dividing by $N$ and letting $N\to\infty$, any subsequential limit $y$ of $\{y^{(N)}\}$ satisfies
$\sum_{i=1}^{K}\tilde{a}_{\tilde{t},i}y_{i}\leq\pi_{\tilde{t}},\qquad\forall\tilde{t}\in\widetilde{\mathbb{T}}$,
so $y$ is feasible for (S2) and $U^{\mathrm{\scriptscriptstyle T}}y\leq\rho$.
The objective value on $G_{N}$ is
$\mathrm{Opt}(G_{N})=\sum_{i=1}^{K}u_{i}Z^{(N)}_{i}=NU^{\mathrm{\scriptscriptstyle T}}y^{(N)},$
so
$\limsup_{N\to\infty}\frac{\mathrm{Opt}(G_{N})}{N}=\limsup_{N\to\infty}U^{\mathrm{\scriptscriptstyle T}}y^{(N)}\leq\rho.$
Step 3B: Lower bound. Define integer vectors
$Z^{(N)}:=\bigl(\lfloor Nz^{\star}_{1}\rfloor,\dots,\lfloor Nz^{\star}_{K}\rfloor\bigr),\qquad y^{(N)}:=\frac{1}{N}Z^{(N)}.$
Then $y^{(N)}\to z^{\star}$, and for each coarse type $\tilde{t}$,
$\sum_{i=1}^{K}\tilde{a}_{\tilde{t},i}y^{(N)}_{i}\leq\sum_{i=1}^{K}\tilde{a}_{\tilde{t},i}z^{\star}_{i}\leq\pi_{\tilde{t}}$,
so $\sum_{i}\tilde{a}_{\tilde{t},i}Z_{i}^{(N)}\leq N\pi_{\tilde{t}}+O(1)$. Together with (S1), this implies $\sum_{i}\tilde{a}_{\tilde{t},i}Z_{i}^{(N)}\leq n_{\tilde{t}}(N)$ for all large $N$, so the coarse-type supply of vertices suffices to realize $Z^{(N)}$ templates.
To realize these templates, fix $i$. Label the vertices of $H_{i}$ as $w_{1},\dots,w_{s_{i}}$ with coarse types $\tilde{t}_{1},\dots,\tilde{t}_{s_{i}}$. For each $(i,\ell)$ choose a subset
$A^{(i)}_{\ell}(N)\subseteq\{v\in V_{N}:\tilde{\tau}(v)=\tilde{t}_{\ell}\}$
of size $Z^{(N)}_{i}$, such that all sets $A^{(i)}_{\ell}(N)$ are disjoint across all $(i,\ell)$; this is possible by the previous capacity inequality.
For fixed $i$, consider the induced $s_{i}$-partite undirected graph on parts $A^{(i)}_{1}(N),\dots,A^{(i)}_{s_{i}}(N)$. Let $L_{i}$ be the number of directed edges in $H_{i}$. If $u\in A^{(i)}_{\ell}(N)$ and $v\in A^{(i)}_{m}(N)$ with $\ell\neq m$, we connect $u$ and $v$ by an undirected edge if all directed edges between $w_{\ell}$ and $w_{m}$ required by $H_{i}$ (and medically admissible) are present between $u$ and $v$ in $G_{N}$. Each such undirected edge appears with probability at least $p_{0}^{L_{i}}=:p_{i}>0$, and different pairs $(u,v)$ use disjoint sets of underlying directed edges, so these edges are independent.
Thus, for each $i$, the $s_{i}$-partite graph on $A^{(i)}_{1}(N),\dots,A^{(i)}_{s_{i}}(N)$ satisfies the hypotheses of Lemma S1 with $k=s_{i}$, $n=Z^{(N)}_{i}$, and parameter $p_{i}>0$. If $z^{\star}_{i}>0$, then $Z^{(N)}_{i}\to\infty$ and Lemma S1 implies that this graph contains $Z^{(N)}_{i}$ vertex-disjoint complete $s_{i}$-partite subgraphs with probability 1o(1)1-o(1). Each such subgraph corresponds to one copy of $H_{i}$. If $z^{\star}_{i}=0$, then $Z^{(N)}_{i}=O(1)$ and failure to realize these copies changes the total utility by at most $O(1)=o(N)$.
Since there are finitely many templates, a union bound shows that with probability 1o(1)1-o(1) we can realize all templates $H_{i}$ with counts $Z^{(N)}_{i}$ simultaneously. On this event, we obtain an exchange plan with template counts $Z^{(N)}$ and utility
$\sum_{i=1}^{K}Z^{(N)}_{i}u_{i}=N\,U^{\mathrm{\scriptscriptstyle T}}z^{\star}+O(1)=\rho N+o(N).$
Hence $\liminf_{N\to\infty}\frac{\mathrm{Opt}(G_{N})}{N}\geq\rho$, proving (S3).
Step 4: Splitting coarse usage into fine types and encoding priority. We now refine coarse-type usage to fine types (including the protected feature) in a way that captures the desired subgroup prioritization.
Step 4A: Coarse usage per type. For each coarse type $\tilde{t}$, define
$\lambda_{\tilde{t}}:=\sum_{i=1}^{K}\tilde{a}_{\tilde{t},i}z^{\star}_{i}\leq\pi_{\tilde{t}}$.
Step 4B: Splitting into fine-type usages. Choose nonnegative numbers $\lambda_{t}$ for $t=(\tilde{t},a)\in\mathbb{T}$ such that
$\lambda_{(\tilde{t},0)}+\lambda_{(\tilde{t},1)}=\lambda_{\tilde{t}},\qquad\lambda_{t}\leq\pi_{t}\quad\forall t.$
For each prioritized quadruple $(b_{1},b_{2},i,j)\in\mathcal{P}$ with coarse type $\tilde{t}=(b_{1},b_{2},r_{j})$ and fine types
$t^{(\mathrm{prio})}=(\tilde{t},i),\qquad t^{(\mathrm{unprio})}=(\tilde{t},1-i),$
we further require that either $\lambda_{t^{(\mathrm{unprio})}}=0$ or $\lambda_{t^{(\mathrm{prio})}}=\pi_{t^{(\mathrm{prio})}}$. This is achieved by the rule
$(\lambda_{t^{(\mathrm{prio})}},\lambda_{t^{(\mathrm{unprio})}})=\begin{cases}(\lambda_{\tilde{t}},0),&\lambda_{\tilde{t}}\leq\pi_{t^{(\mathrm{prio})}},\\[1.99997pt] (\pi_{t^{(\mathrm{prio})}},\lambda_{\tilde{t}}-\pi_{t^{(\mathrm{prio})}}),&\lambda_{\tilde{t}}>\pi_{t^{(\mathrm{prio})}}.\end{cases}$
For coarse types not in any prioritized pair, we pick any feasible split.
Step 4C: Fine-type template counts. We refine the coarse counts $\tilde{a}_{\tilde{t},i}$ into $b_{t,i}\geq 0$ so that
$\sum_{a=0}^{1}b_{(\tilde{t},a),i}=\tilde{a}_{\tilde{t},i}\quad\forall\tilde{t},i,\qquad\sum_{i=1}^{K}b_{t,i}z^{\star}_{i}=\lambda_{t}\quad\forall t\in\mathbb{T}.$ (S4)
If $\lambda_{\tilde{t}}>0$, we set
$b_{(\tilde{t},a),i}:=\tilde{a}_{\tilde{t},i}\,\frac{\lambda_{(\tilde{t},a)}}{\lambda_{\tilde{t}}},\quad a\in\{0,1\},$
and if $\lambda_{\tilde{t}}=0$, we set $b_{(\tilde{t},a),i}=0$. A direct check shows that (S4) holds.
Step 5: Constructing matching fine-type counts in $G_{N}$. We now pass from the limiting fractions $\lambda_{t}$ and $b_{t,i}$ to integer counts in $G_{N}$.
Step 5A: Target number of matched vertices per fine type. For each fine type $t$ and $N$, let
$M^{(N)}_{t}:=\bigl\lfloor N\lambda_{t}\bigr\rfloor.$
By (S1) and $\lambda_{t}\leq\pi_{t}$, we have $M^{(N)}_{t}\leq n_{t}(N)$ for all large $N$, so we can choose
$V^{\mathrm{use}}_{t}(N)\subseteq\{v\in V_{N}:\tau(v)=t\},\qquad|V^{\mathrm{use}}_{t}(N)|=M^{(N)}_{t}$.
For each prioritized quadruple $(b_{1},b_{2},i,j)\in\mathcal{P}$ with fine types $t^{(\mathrm{prio})},t^{(\mathrm{unprio})}$, our construction of $\lambda_{t}$ implies that, up to $o(N)$ vertices, either only the prioritized type is used or the prioritized type is essentially saturated. This exactly matches the asymptotic priority requirement in the definition of $\mathrm{Opt}^{\mathrm{prio}}_{\varepsilon_{N}}(G_{N})$ for some sequence $\varepsilon_{N}\downarrow 0$.
Step 5B: Matching fine-type/template counts and realization. Let $\gamma_{t,i}:=b_{t,i}z^{\star}_{i}$. We seek integers $m^{(N)}_{t,i}$ such that
$\sum_{i=1}^{K}m^{(N)}_{t,i}=M^{(N)}_{t}\quad\forall t\in\mathbb{T},\qquad\sum_{t\in\mathbb{T}}m^{(N)}_{t,i}=s_{i}Z^{(N)}_{i}\quad\forall i,$
and $m^{(N)}_{t,i}=N\gamma_{t,i}+O(1)$ as $N\to\infty$. This can be obtained by a standard rounding argument on the finite transportation polytope: start from $m^{(N)}_{t,i}:=\lfloor N\gamma_{t,i}\rfloor$ and adjust finitely many entries by $O(1)$ to match the row and column sums exactly. The total adjustment per $(t,i)$ is $O(1)$, so this changes the utility by at most $O(1)=o(N)$.
For each pair $(t,i)$, partition $V^{\mathrm{use}}_{t}(N)$ into disjoint subsets of size $m^{(N)}_{t,i}$ and assign each subset to the appropriate position in the $Z^{(N)}_{i}$ copies of $H_{i}$. This yields, for each template $i$, parts $A^{(i)}_{1}(N),\dots,A^{(i)}_{s_{i}}(N)$ of size $Z^{(N)}_{i}$ with prescribed fine (and hence coarse) types.
We now apply exactly the same Lemma S1-based construction as in Step 3B, restricted to these parts. The independence and lower bounds on edge probabilities are unchanged, so with probability 1o(1)1-o(1) we can realize the desired $Z^{(N)}_{i}$ disjoint copies of $H_{i}$ using precisely these vertices, for all $i$ with $z^{\star}_{i}>0$. Templates with $z^{\star}_{i}=0$ contribute $O(1)$ vertices and can be handled arbitrarily, with $o(N)$ effect.
Thus, with probability 1o(1)1-o(1), for all large $N$ there exists an exchange plan that:
  • uses exactly the vertices in $\bigcup_{t}V^{\mathrm{use}}_{t}(N)$ up to $O(1)$,
  • has template counts $Z^{(N)}$, and
  • satisfies the asymptotic priority requirement with sequence $\varepsilon_{N}$.
The corresponding utility is
$\sum_{i=1}^{K}Z^{(N)}_{i}u_{i}+O(1)=\rho N+o(N),$
so almost surely
$\mathrm{Opt}^{\mathrm{prio}}_{\varepsilon_{N}}(G_{N})\geq\rho N+o(N).$
Step 6: Conclusion. From Step 3 we know $\mathrm{Opt}(G_{N})=\rho N+o(N)$ almost surely, and from Step 5 we know $\mathrm{Opt}^{\mathrm{prio}}_{\varepsilon_{N}}(G_{N})\geq\rho N+o(N)$ almost surely. Hence
$\mathrm{Opt}(G_{N})-\mathrm{Opt}^{\mathrm{prio}}_{\varepsilon_{N}}(G_{N})=o(N)\quad\text{almost surely,}$
proving the proposition. ∎
@@
We bound the price of fairness in terms of the number of matched vertices and show it is at most the explicit quantity $L$ in the statement. The argument is purely type-level and ignores graph structure until the final step.
Step 1: Local loss for a fixed $(b_{1},b_{2},r)$. Fix $(b_{1},b_{2})$ and $r$, and consider the two fine types
$(b_{1},b_{2},r,1)\quad\text{and}\quad(b_{1},b_{2},r,0),$
with proportions $\mu_{b_{1},b_{2},r,1}$ and $\mu_{b_{1},b_{2},r,0}$. Let
$\overline{\mu}_{r,a}:=\sum_{b_{1},b_{2}}\mu_{b_{1},b_{2},r,a},\qquad a\in\{0,1\},$
be the total mass at level $r$ in group $A=a$.
In the unconstrained optimum, the total fraction of matched vertices of these two types can be written as
$p\,(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})$
for some $p\in[0,1]$, which we can think of as the local selection probability at $(b_{1},b_{2},r)$. We now compare this with an allocation that enforces the stronger requirement that the selection probabilities at level $r$ for $A=0$ and $A=1$ are exactly equal:
$\frac{\text{matches in }(A=1,R=r)}{\overline{\mu}_{r,1}}=\frac{\text{matches in }(A=0,R=r)}{\overline{\mu}_{r,0}}.$
This is stricter than the original fairness constraint, so any loss bound obtained here is conservative.
If the ratio $\mu_{b_{1},b_{2},r,1}/\mu_{b_{1},b_{2},r,0}$ exceeds the global ratio $\overline{\mu}_{r,1}/\overline{\mu}_{r,0}$, then this type is “too rich” in $A=1$ individuals. A straightforward two-type calculation (keeping all possible $A=0$ matches and enforcing the global ratio) shows that the fair allocation must lose at most
$\frac{\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}-\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\,\overline{\mu}_{r,0}}$
of the local matched mass. When the inequality is reversed, the type is “too rich” in $A=0$, and by symmetry (exchanging the roles of $A=0$ and $A=1$) the maximal local loss is
$\frac{\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}-\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\,\overline{\mu}_{r,1}}.$
Taking the larger of these two expressions gives the loss bound for fixed $(b_{1},b_{2},r)$.
Step 2: Global bound across all $(b_{1},b_{2},r)$. At a fixed sensitization level $r$, the total matches in the unconstrained optimum are a sum over $(b_{1},b_{2})$ of $M^{*}_{b_{1},b_{2},r}(N)$, and the number of matches in any fair allocation is the corresponding sum after trimming. Hence, at level $r$, the (asymptotic) fractional loss is a convex combination of the local losses $\mathrm{Loss}_{b_{1},b_{2},r}$, and so is bounded by their maximum over $(b_{1},b_{2})$.
Taking also the maximum over $r$ yields the loss bound in the statement.
Step 3: Implementability in $G_{N}$ via subgroup prioritization. We now argue that the type-level bound $L$ governs the asymptotic price of fairness in the random graphs $G_{N}$.
For each triple $(b_{1},b_{2},r)$, the trimming operation in Step 1 can be implemented by prioritizing between the fine types $(b_{1},b_{2},r,1)$ and $(b_{1},b_{2},r,0)$, as in Section 4.1: we decide, for each $(b_{1},b_{2},r)$, whether to prioritize $A=1$ or $A=0$, and collect these choices into a priority set $\mathcal{P}$.
Consider the constrained optimization problem with subgroup prioritization constraints (6)–(7) associated with $\mathcal{P}$. By Proposition 1, there exists a sequence $\varepsilon_{N}\downarrow 0$ such that we can impose the asymptotic priority requirements for all $(b_{1},b_{2},i,j)\in\mathcal{P}$ while losing only $o(N)$ matches relative to the unconstrained optimum.
Within this prioritized class, we can then discard matches exactly as in the type-level construction of Step 1 to enforce equality of selection probabilities between $A=0$ and $A=1$ at each $r$. The additional loss caused by this trimming is, by Step 2, at most a fraction $L$ of the total number of matches, that is, at most $L\,\mathrm{Opt}(G_{N})+o(N)$ matches in $G_{N}$.
Thus, almost surely, there exists an allocation satisfying the stronger condition of equal selection probabilities at each $r$ whose total number of matches is at least $(1-L)\,\mathrm{Opt}(G_{N})-o(N)$. Since the fairness constraint (4) only requires the difference in selection probabilities at each $r_{j}$ to be bounded by $l_{j}$ and not necessarily equal to zero, any optimal solution under (4) is at least as efficient as this equalized allocation. Consequently, the asymptotic price of fairness due to (4) is no greater than $L$. ∎